OpenStudy (anonymous):

How many solutions does this system have? -3x+6y=10 -3x+6y=-4

OpenStudy (anonymous):

If you had a system with x + y = 5 and x + y = 6, what would be your first thought? There is no pair of numbers such that when we add them together, they make two different numbers - every sum of a pair of numbers is unique. Hopefully that notion should help with this question.

OpenStudy (anonymous):

none

OpenStudy (anonymous):

these are two parallel lines, so no solution

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (anonymous):

Can you help me with one more?

OpenStudy (anonymous):

Which method would be the simplest way to solve the system? 7x+5y=19 -7x-2y=-16

OpenStudy (anonymous):

Simultaneous equations, adding the second row to the first to cancel out the x's

OpenStudy (anonymous):

So basicly elimination?

OpenStudy (anonymous):

Aye, that one :) (sorry, never could remember the names of sim equation methods!)

OpenStudy (anonymous):

It's totally fine, you helped me still. Thanks!!!

OpenStudy (anonymous):

No problem :)